OpenStudy (shelby1290):

Use the suggested method to sketch the graph of each linear relationship. a) y= -4x+3 (slope, y-intercept method) b) 2x-y=4 (x-,y-intercept method) c) 3x=6y-3 (table of values)

OpenStudy (texaschic101):

a.) y = -4x + 3 --- slope, y int method.. In y = mx + b form, the slope is in the m position and the y intercept is in the b position. So...the slope is -4 (same as -4/1) and the y int is 3, or (0,3). First we plot out y int points (0,3).....and since our slope is -4/1, we come down 4 units (because of the negative), and go right 1 unit, then down 4 units, the to the right 1.. If you want to make sure you are doing it right, find the x int by subbing in 0 for y. y = -4x + 3 0 = -4x + 3 4x = 3 x = 3/4....so the x int is (3/4,0) And following the slope, coming down from (0,3).....coming down 4 and to the right 1, you should cross the x axis at (3/4,0) Looks like this .. http://www.meta-calculator.com/online/?panel-102-graph&data-bounds-xMin=-10&data-bounds-xMax=10&data-bounds-yMin=-7.28&data-bounds-yMax=7.28&data-equations-0=%22y%20%3D%20-4*x%20%2B%203%22&data-rand=undefined&data-hideGrid=false any questions about this one ?

OpenStudy (texaschic101):

b.) 2x - y = 4 ....x and y intercept method To find x intercept, sub in 0 for y 2x - 0 = 4 2x = 4 x = 2.....so the x intercept is (2,0) To find the y intercept, sub in 0 for x 2(0) - y = 4 -y = 4 y = -4 ...so the y intercept is (0,-4) so now you plot your points (2,0) and (0,-4)......connect them with a solid line Looks like this... http://www.meta-calculator.com/online/?panel-102-graph&data-bounds-xMin=-10&data-bounds-xMax=10&data-bounds-yMin=-7.28&data-bounds-yMax=7.28&data-equations-0=%222*x%20-%20y%20%3D%204%22&data-rand=undefined&data-hideGrid=false

OpenStudy (texaschic101):

c.) 3x - 6y = 3.....now we will sub in numbers for x to find y Lets let x = 0 3(0) - 6y = 3 -6y = 3 y = -1/2 Lets let x be 1 3(1) - 6y = 3 3 - 6y = 3 -6y = 3 - 3 -6y = 0 y = 0 Lets let x be 2 3(2) - 6y = 3 6 - 6y = 3 -6y = 3 - 6 -6y = -3 y = 1/2 one more...let x be 3 3(3) - 6y = 3 9 - 6y = 3 -6y = 3 - 9 -6y = -6 y = 1 so...out table now has 4 sets of points... x | y ---------------- 0 -1/2 1 0 2 1/2 3 1 Now it is just a matter of plotting your points (0,-1/2) , (1,0), (2,1/2) , (3,1) Looks like this... http://www.meta-calculator.com/online/?panel-102-graph&data-bounds-xMin=-10&data-bounds-xMax=10&data-bounds-yMin=-7.28&data-bounds-yMax=7.28&data-equations-0=%223*x%20-%206*y%20%3D%203%22&data-rand=undefined&data-hideGrid=false

OpenStudy (shelby1290):

@texaschic101 wow thank you so much